# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu3,11*, Wenhao Huang2,3

### v2

@@ Abstract
In this paper, we introduce MIO, a novel foundation model built on multimodal tokens, capable of understanding and generating speech, text, images, and videos in an end-to-end, autoregressive manner. While the emergence of large language models (LLMs) and multimodal large language models (MM-LLMs) propels advancements in artificial general intelligence through their versatile capabilities, they still lack true any-to-any understanding and generation. Recently, the release of GPT-4o has showcased the remarkable potential of any-to-any LLMs for complex real-world tasks, enabling omnidirectional input and output across images, speech, and text. However, it is closed-source and does not support the generation of multimodal interleaved sequences. To address this gap, we present MIO, which is trained on a mixture of discrete tokens across four modalities using causal multimodal modeling. MIO undergoes a four-stage training process: (1) alignment pre-training, (2) interleaved pre-training, (3) speech-enhanced pre-training, and (4) comprehensive supervised fine-tuning on diverse textual, visual, and speech tasks. Our experimental results indicate that MIO exhibits competitive, and in some cases superior, performance compared to previous dual-modal baselines, any-to-any model baselines, and even modality-specific baselines. Moreover, MIO demonstrates advanced capabilities inherent to its any-to-any feature, such as interleaved video-text generation, chain-of-visual-thought reasoning, visual guideline generation, instructional image editing, etc.
@@ 1 Introduction
The advent of Large Language Models (LLMs) is commonly considered the dawn of artificial general intelligence (AGI) [1, 2], given their generalist capabilities such as complex reasoning [3], role playing [4], and creative writing [5]. However, original LLMs lack multimodal understanding capabilities. Consequently, numerous multimodal LLMs (MM-LLMs) have been proposed, allowing LLMs to understand images [6, 7], audio [8, 9, 10, 11], and other modalities [12, 13, 14]. These MM-LLMs typically involve an external multimodal encoder, such as EVA-CLIP [15] or CLAP [16], with an alignment module such as Q-Former [6] or MLP [17] for multimodal understanding. These modules align non-textual-modality data features into the embedding space of the LLM backbone.
@@
Models Emu1 [18] Emu2 [19] SEED-LLaMA [20] AnyGPT [21] CM3Leon [22], Chameleon [23] Gemini [24] Transfusion [25] MIO (ours)
@@
Speech Input/Output / / / / / / /
Video Input/Output / / / / / / /
@@
Another line of work involves building any-to-any and end-to-end MM-LLMs that can input and output non-textual modality data. Typically, there are four approaches: (1) Discrete-In-Discrete-Out (DIDO): Non-textual modality data is discretized using vector quantization techniques [26, 27] and then fed into LLMs [20, 21, 28]. (2) Continuous-In-Discrete-Out (CIDO): The LLM backbones intake densely encoded non-textual modality data features and generate their quantized representations [29, 30]. (3) Continuous-In-Continuous-Out (CICO): The LLMs both understand and generate non-textual modality data in their densely encoded representations [18, 19, 31, 32, 33]. (4) Autoregression + Diffusion (AR + Diff): The autoregressive and diffusion modeling are integrated in a unified LLM [25, 34, 35]. Although these works have succeeded in building MM-LLMs unifying understanding and generation, they exhibit some drawbacks, as illustrated in Table 1. For example, Emu1 [18] and Emu2 [19] explore the autoregressive modeling of three modalities: text, images, and videos. SEED-LLaMA [20] proposes a new image quantizer aligned with LLMs’ embedding space and trains the MM-LLMs on images and videos. However, neither considers the speech modality, which is heterogeneous from visual modalities like videos and images. Although AnyGPT [21] has explored settings involving four modalities, including text, image, speech, and music, it lacks video-related abilities, voice synthesis, and comprehensive multi-task supervised fine-tuning, leading to limited multimodal instruction-following and reasoning capabilities. Furthermore, AR + Diff approaches, such as Transfusion [25], suffer from limited multimodal understanding capabilities because the multimodal inputs are noised for denoising modeling, and the image tokenizer used is suitable for generation rather than understanding.
@@
Moreover, most of current MM-LLMs are typically dual-modal, combining text with another modality, such as images. Although previous works, such as Meta-Transformer [13] and Unified-IO 2 [36], have explored omni-multimodal understanding settings with more than two non-textual modalities, they still lag significantly behind their dual-modal counterparts, especially in terms of multimodal instruction-following capabilities. Moreover, these MM-LLMs are typically focused on understanding only, neglecting the important aspect of multimodal generation. Several works have enabled LLMs to call external tools to address this issue. For example, HuggingGPT [37] generates textual image descriptions for external diffusion models to synthesize images. GPT-4 [1] can utilize either an image generator like DALL-E 3 [38] or a text-to-speech (TTS) tool like Whisper [39] to support multimodal generation. [^1] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/ However, these methods are not end-to-end, relying on the text modality as an interface.
@@
Recently, the release of GPT-4o has demonstrated the capabilities of any-to-any and end-to-end foundation models. [^2] https://openai.com/index/hello-gpt-4o/ It is the first foundational model to accept multimodal tokens as inputs and generate multimodal tokens within a unified model while also demonstrating strong abilities in complex multimodal instruction-following, reasoning, planning, and other generalist capabilities. Furthermore, as the continuous scaling up of LLMs in the community depletes high-quality language tokens, GPT-4o verifies a new source of data for LLM training: multimodal tokens. This approach suggests that the next generation AGI could derive more knowledge from multimodal tokens when language tokens are exhausted. However, GPT-4o is closed source and focuses primarily on end-to-end support for speech I/O, image I/O, 3D generation, and video understanding. Its recent open-source “alternatives”, such as VITA [40], still lack the ability to generate data of all supported modalities, particularly for the generation of multimodal interleaved sequences.
@@
To address the aforementioned issues, we introduce MIO (Multimodal Input and Output, or Multimodal Interleaved Output), the first open-source any-to-any foundation model that unifies multimodal understanding and generation across four modalities–text, image, speech (with voice), and video, while enabling the generation of multimodal interleaved sequences. Specifically, MIO is built on discrete multimodal tokens that capture both semantic representations through contrastive loss and low-level features via reconstruction loss [41, 42] from raw multimodal data. Due to the consistent data format shared with textual corpora, the model can treat non-textual modalities as “foreign languages”, allowing it to be trained with the next-token-prediction paradigm. Note that since the representation of an image remains the same whether it is used as an input or an output, our model flexibly supports multimodal interleaved sequence generation, where an image functions simultaneously for both understanding and generation. Moreover, we employ a three-stage pre-training process, along with an additional SFT stage, to effectively train the model for modality expansion.
@@
Our experimental results show that MIO, trained on a mixture of four modalities, demonstrates competitive performance compared to its dual-modal counterparts and previous any-to-any multimodal language model baselines. Additionally, MIO is the first model to demonstrate interleaved video-text generation, chain-of-visual-thought reasoning, and other emergent abilities relying on any-to-any and multimodal interleaved output features (c.f.3.4).
@@ 2.1 Modeling
In our work, we use SEED-Tokenizer [41] as our image tokenizer and SpeechTokenizer [42] as our speech tokenizer. SEED-Tokenizer encodes images using a ViT [43] derived from BLIP-2 [6], and then converts the encoded features into fewer tokens with causal semantics via Q-Former [6]. These features are subsequently quantized into discrete tokens that are well-aligned with the language model backbone’s textual space. The codebook size for these discrete image tokens is 8192. SEED-Tokenizer transforms each image into a 224x224 resolution and quantizes it into 32 tokens. We use two special tokens, <image> and </image>, to indicate the start and end of the image tokens per image, respectively.
@@
In terms of speech, SpeechTokenizer [42] leverages an 8-layer RVQ [44] to tokenize speech into tokens with 8 codebooks, with each codebook derived from one layer. Since the first layer’s quantization output is distilled from HuBERT [45], which encodes more semantic information, SpeechTokenizer can separate content tokens and timbre tokens from a quantized speech. The first-layer quantization is treated as content quantization, while the remaining layers’ quantization is treated as timbre quantization. SpeechTokenizer encodes speech into 50 tokens per second for each codebook, resulting in 400 tokens per second with all eight codebooks. To improve context efficiency, we drop the last four layers’ codebooks and only use the content codebook and the first three timbre codebooks. Our vocabulary size for the speech modality is $1024\times 4=4096$.
Since the open-source pretraining-level speech data is collected from individuals with diverse voices, the timbre tokens exhibit a relatively random and noisy pattern, while the content tokens are more fixed-pattern and better aligned with the corresponding transcriptions. Given these priors in speech tokens, it is important to choose the proper interleaving mode of speech tokens [46]. We denote the four codebooks as $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$, where $\mathcal{A}$ is the codebook for content tokens and the remaining three are for timbre tokens. For simplicity, assuming that we have only two tokens for each codebook in a tokenized speech sequence (i.e., $a_{1}a_{2}$, $b_{1}b_{2}$, $c_{1}c_{2}$, and $d_{1}d_{2}$), there are two interleaving patterns for causal multimodal modeling: (1) sequential interleaving pattern: $a_{1}a_{2}b_{1}b_{2}c_{1}c_{2}d_{1}d_{2}$ and (2) alternating interleaving pattern: $a_{1}b_{1}c_{1}d_{1}a_{2}b_{2}c_{2}d_{2}$.
@@
As illustrated in Figure 1, the speech and images, including video frames, are tokenized by SpeechTokenizer [42] and SEED-Tokenizer [41], respectively. We add the 4096 speech tokens and 8192 image tokens to the LLM’s vocabulary. In addition, we introduce four new special tokens, namely <image>, </image>, <spch>, and </spch>, to the vocabulary. Consequently, the embedding layer of the LLM backbone and the language modeling head are extended by $4096+8192+4=12292$ to support the embedding and generation of these new tokens. The image tokens contain causal semantics due to the use of a Causal Q-Former [41], and the speech tokens are intrinsically causal due to their temporal nature. Therefore, these multimodal tokens are as suitable for autoregressive training as textual tokens, allowing us to unify the training objectives for understanding and generation of multimodal tokens into next-token-prediction with cross-entropy loss. The training objective is thus:
@@
where $x_{t}$ represents the discrete multimodal tokens, and $\theta$ denotes the parameters of the LLM backbone. We use the pre-trained model, Yi-6B-Base [47], for initialization.
Furthermore, to eliminate the computational inefficiency caused by <pad> tokens, we use the masked packing strategy [36, 28, 48]. Specifically, the samples are concatenated along the sequence length dimension until the context window is full. Then, we construct the causal attention mask for the tokens of each sample and mask out all the tokens of the other samples.
@@
After the generation of multimodal tokens, it is essential to use modality-specific decoders to reconstruct the images or speech from the codes. Specifically, for image tokens, we directly utilize SEED-Tokenizer’s decoder, which involves an MLP projection to convert the discrete codes into dense latents. These latents condition an off-the-shelf diffusion model [49] to generate the images in the pixel space [41]. The vanilla SpeechTokenizer [42] involves generating timbre tokens through a non-autoregressive model outside the language model, and then feeding the concatenated content and timbre tokens into the SpeechTokenizer decoder to synthesize speech. In our work, to inject the timbre priors into the multimodal language model itself, the timbre tokens are also generated by the autoregressive language model.
@@ 2.2 Pre-Training
As shown in Table 2, we use a three-stage strategy for pre-training, with each stage targeting different objectives. The three stages are: (1) Alignment Pre-training: This stage focuses on learning a multimodal representation more aligned with the language space. (2) Interleaved Pre-training: This stage aims to obtain a multimodal representation with richer contextual semantics. (3) Speech-enhanced Pre-training: This stage specifically enhances the model’s speech-related capabilities, while concurrently replaying data from other modalities. For more details on the pre-training data and its processing procedures, we refer the readers to Appendix B.
Table 2: Pre-training stages and their details. We use “Inter” to denote “Interleaved” for short. We provide batch sizes for each data type per GPU in image-text pair data:language-only data:(image-text interleaved data + video data):speech-text pair data. See Appendix B and Appendix C for more details including pre-training data sources, data cleaning procedures, pre-training hyperparameters, etc.
@@
In this stage, we extend the data used for pre-training to include image-text interleaved data (including video-text data) as a novel image-text dependency pattern. The image-text interleaving pattern has a different nature compared to pairing patterns. Although [6] and [18] argued that interleaved image-text data mainly serves for multimodal in-context learning, we argue that it is also essential for context-aware image generation where images are generated based on specific context, rather than a precise description of the image content. For example, in image-text interleaved data, the text might serve as the image’s preceding or continuing context, rather than its description. This pattern significantly differs from the previous descriptive image generation demonstrated in image-text paired data, where images are generated based on precise and detailed text that clearly describe the content of the images [30]. Therefore, context-aware image generation is essential for tasks such as chain-of-visual-thought reasoning or visual storytelling [30, 50], where images are generated without textual descriptions. Due to the lack of benchmarks and evaluation metrics for context-aware image generation, we provide some demonstrations in §3.4 to showcase the potential of our model in visual storytelling, interleaved video-text generation, instructional image editing, chain-of-visual-thought reasoning, multimodal in-context learning, etc.
@@
We refer the reader to Appendix C for more details about the hyperparameters and prompt templates used for pre-training.

### v3

# Zhaoxiang Zhang9, Yuanxing Zhang10, Ge Zhang3,7, Ke Xu1, Jie Fu11*, Wenhao Huang2,3

@@ Abstract
In this paper, we introduce MIO, a novel foundation model built on multimodal tokens, capable of understanding and generating speech, text, images, and videos in an end-to-end, autoregressive manner. While the emergence of large language models (LLMs) and multimodal large language models (MM-LLMs) propels advancements in artificial general intelligence through their versatile capabilities, they still lack true any-to-any understanding and generation. Recently, the release of GPT-4o has showcased the remarkable potential of any-to-any LLMs for complex real-world tasks, enabling omnidirectional input and output across images, speech, and text. However, it is closed-source and does not support the generation of multimodal interleaved sequences. To address this gap, we present MIO, which is trained on a mixture of discrete tokens across four modalities using causal multimodal modeling. MIO undergoes a four-stage training process: (1) alignment pre-training, (2) interleaved pre-training, (3) speech-enhanced pre-training, and (4) comprehensive supervised fine-tuning on diverse textual, visual, and speech tasks. Our experimental results indicate that MIO exhibits competitive, and in some cases superior, performance compared to previous dual-modal baselines, any-to-any model baselines, and even modality-specific baselines. Moreover, MIO demonstrates advanced capabilities inherent to its any-to-any feature, such as interleaved video-text generation, chain-of-visual-thought reasoning, visual guideline generation, instructional image editing, etc. Codes and models are available at https://github.com/MIO-Team/MIO.
@@ 1 Introduction
The advent of Large Language Models (LLMs) is commonly considered the dawn of artificial general intelligence (AGI) (OpenAI et al., 2023; Bubeck et al., 2023), given their generalist capabilities such as complex reasoning (Wei et al., 2022), role playing (Wang et al., 2023c), and creative writing (Wang et al., 2024a). However, original LLMs lack multimodal understanding capabilities. Consequently, numerous multimodal LLMs (MM-LLMs) have been proposed, allowing LLMs to understand images (Li et al., 2023b; Alayrac et al., 2022), audio (Borsos et al., 2023; Rubenstein et al., 2023; Tang et al., 2023; Das et al., 2024), and other modalities (Lyu et al., 2023; Zhang et al., 2023d; Moon et al., 2023). These MM-LLMs typically involve an external multimodal encoder, such as EVA-CLIP (Sun et al., 2023b) or CLAP (Elizalde et al., 2022), with an alignment module such as Q-Former (Li et al., 2023b) or MLP (Liu et al., 2023b) for multimodal understanding. These modules align non-textual-modality data features into the embedding space of the LLM backbone.
@@
Models Emu1
 (Sun et al., 2023c)
Emu2
 (Sun et al., 2023a)
SEED-LLaMA
 (Ge et al., 2023b)
AnyGPT
 (Zhan et al., 2024)
CM3Leon
 (Yu et al., 2023), Chameleon
 (Team, 2024)
Gemini
 (Reid et al., 2024)
Transfusion
 (Zhou et al., 2024)
MIO
@@
Speech I/O / / / / / / /
Video I/O / / / / / / /
@@
Another line of work involves building any-to-any and end-to-end MM-LLMs that can input and output non-textual modality data. Typically, there are four approaches: (1) Discrete-In-Discrete-Out (DIDO): Non-textual modality data is discretized using vector quantization techniques (van den Oord et al., 2017; Esser et al., 2020) and then fed into LLMs (Ge et al., 2023b; Zhan et al., 2024; Liu et al., 2024). (2) Continuous-In-Discrete-Out (CIDO): The LLM backbones intake densely encoded non-textual modality data features and generate their quantized representations (Diao et al., 2023; Team et al., 2023). (3) Continuous-In-Continuous-Out (CICO): The LLMs both understand and generate non-textual modality data in their densely encoded representations (Sun et al., 2023c; a; Dong et al., 2023; Zheng et al., 2023; Wu et al., 2023). (4) Autoregression + Diffusion (AR + Diff): The autoregressive and diffusion modeling are integrated in a unified LLM (Zhou et al., 2024; Xie et al., 2024; Li et al., 2024b). Although these works have succeeded in building MM-LLMs unifying understanding and generation, they exhibit some drawbacks, as illustrated in Table 1. For example, Emu1 (Sun et al., 2023c) and Emu2 (Sun et al., 2023a) explore the autoregressive modeling of three modalities: text, images, and videos. SEED-LLaMA (Ge et al., 2023b) proposes a new image quantizer aligned with LLMs’ embedding space and trains the MM-LLMs on images and videos. However, neither considers the speech modality, which is heterogeneous from visual modalities like videos and images. Although AnyGPT (Zhan et al., 2024) has explored settings involving four modalities, including text, image, speech, and music, it lacks video-related abilities, voice synthesis, and comprehensive multi-task supervised fine-tuning, leading to limited multimodal instruction-following and reasoning capabilities. Furthermore, AR + Diff approaches, such as Transfusion (Zhou et al., 2024), suffer from limited multimodal understanding capabilities because the multimodal inputs are noised for denoising modeling, and the image tokenizer used (i.e., VAE (Kingma & Welling, 2013)) is suitable for image generation rather than image understanding.
@@
Moreover, most of current MM-LLMs are typically dual-modal, combining text with another modality, such as images. Although previous works, such as Meta-Transformer (Zhang et al., 2023d) and Unified-IO 2 (Lu et al., 2023), have explored omni-multimodal understanding settings with more than two non-textual modalities, they still lag significantly behind their dual-modal counterparts, especially in terms of multimodal instruction-following capabilities. Moreover, these MM-LLMs are typically focused on understanding only, neglecting the important aspect of multimodal generation. Several works have enabled LLMs to call external tools to address this issue. For example, HuggingGPT (Shen et al., 2023) generates textual image descriptions for external diffusion models to synthesize images. GPT-4 (OpenAI et al., 2023) can utilize either an image generator like DALL-E 3 (Betker et al., 2024) or a text-to-speech (TTS) tool like Whisper (Radford et al., 2022) to support multimodal generation. [^1] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/ However, these methods are not end-to-end, relying on the text modality as an interface.
@@
Recently, the release of GPT-4o has demonstrated the capabilities of any-to-any and end-to-end foundation models. [^2] https://openai.com/index/hello-gpt-4o/ It is the first foundational model to accept multimodal tokens as inputs and generate multimodal tokens within a unified model while also demonstrating strong abilities in complex multimodal instruction-following, reasoning, planning, and other generalist capabilities. Furthermore, as the continuous scaling up of LLMs in the community depletes high-quality language tokens, GPT-4o verifies a new source of data for LLM training: multimodal tokens. This approach suggests that the next generation AGI could derive more knowledge from multimodal tokens when language tokens are exhausted. However, GPT-4o is closed source and focuses primarily on end-to-end support for speech I/O, image I/O, 3D generation, and video understanding. Its recent open-source “alternatives”, such as VITA (Fu et al., 2024), still lack the ability to generate data of all supported modalities, particularly for the generation of multimodal interleaved sequences.
@@
To address the aforementioned issues, we introduce MIO (Multimodal Input and Output, or Multimodal Interleaved Output), the first open-source any-to-any foundation model that unifies multimodal understanding and generation across four modalities–text, image, speech (with voice), and video, while enabling the generation of multimodal interleaved sequences. Specifically, MIO is built on discrete multimodal tokens that capture both semantic representations through contrastive loss and low-level features via reconstruction loss (Ge et al., 2023a; Zhang et al., 2023b) from raw multimodal data. Due to the consistent data format shared with textual corpora, the model can treat non-textual modalities as “foreign languages”, allowing it to be trained with the next-token-prediction. Note that since the representation of an image remains the same whether it is used as an input or an output, our model flexibly supports multimodal interleaved sequence generation, where an image functions simultaneously for both understanding and generation. Moreover, we employ three-stage pre-training with an additional SFT stage to effectively train the model for modality scaling.
@@
Our experimental results show that MIO, trained on a mixture of four modalities, demonstrates competitive performance compared to its dual-modal counterparts and previous any-to-any multimodal language model baselines. Additionally, MIO is the first model to demonstrate interleaved video-text generation, chain-of-visual-thought reasoning, and other emergent abilities relying on any-to-any and multimodal interleaved output features (c.f.3.5).
@@ 2.1 Modeling
In our work, we use SEED-Tokenizer (Ge et al., 2023a) as our image tokenizer and SpeechTokenizer (Zhang et al., 2023b) as our speech tokenizer. SEED-Tokenizer encodes images using a ViT (Dosovitskiy et al., 2021) derived from BLIP-2 (Li et al., 2023b), and then converts the encoded features into fewer tokens with causal semantics via Q-Former (Li et al., 2023b). These features are subsequently quantized into discrete tokens that are well-aligned with the language model backbone’s textual space. The codebook size for these discrete image tokens is 8192. SEED-Tokenizer transforms each image into a 224x224 resolution and quantizes it into 32 tokens. We use two special tokens, <image> and </image>, to indicate the start and end of the image tokens per image, respectively.
@@
In terms of speech, SpeechTokenizer (Zhang et al., 2023b) leverages an 8-layer RVQ (Lee et al., 2022) to tokenize speech into tokens with 8 codebooks, with each codebook derived from one layer. Since the first layer’s quantization output is distilled from HuBERT (Hsu et al., 2021), which encodes more semantic information, SpeechTokenizer can separate content tokens and timbre tokens from a quantized speech. The first-layer quantization is treated as content quantization, while the remaining layers’ quantization is treated as timbre quantization. SpeechTokenizer encodes speech into 50 tokens per second for each codebook, resulting in 400 tokens per second with all eight codebooks. To improve context efficiency, we drop the last four layers’ codebooks and only use the content codebook and the first three timbre codebooks. Our vocabulary size for the speech modality is $1024\times 4=4096$.
Since the open-source pretraining-level speech data is collected from individuals with diverse voices, the timbre tokens exhibit a relatively random and noisy pattern, while the content tokens are more fixed-pattern and better aligned with the corresponding transcriptions. Given these priors in speech tokens, it is important to choose the proper interleaving mode of speech tokens (Copet et al., 2023). We denote the four codebooks as $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$, and $\mathcal{D}$, where $\mathcal{A}$ is the codebook for content tokens and the remaining three are for timbre tokens. For simplicity, assuming that we have only two tokens for each codebook in a tokenized speech sequence (i.e., $a_{1}a_{2}$, $b_{1}b_{2}$, $c_{1}c_{2}$, and $d_{1}d_{2}$), there are two interleaving patterns for causal multimodal modeling: (1) sequential interleaving pattern: $a_{1}a_{2}b_{1}b_{2}c_{1}c_{2}d_{1}d_{2}$ and (2) alternating interleaving pattern: $a_{1}b_{1}c_{1}d_{1}a_{2}b_{2}c_{2}d_{2}$.
@@
As illustrated in Figure 1, the speech and images, including video frames, are tokenized by SpeechTokenizer (Zhang et al., 2023b) and SEED-Tokenizer (Ge et al., 2023a), respectively. We add the 4096 speech tokens and 8192 image tokens to the LLM’s vocabulary. In addition, we introduce four new special tokens, namely <image>, </image>, <spch>, and </spch>, to the vocabulary. Consequently, the embedding layer of the LLM backbone and the language modeling head are extended by $4096+8192+4=12292$ to support the embedding and generation of these new tokens. The image tokens contain causal semantics due to the use of a Causal Q-Former (Ge et al., 2023a), and the speech tokens are intrinsically causal due to their temporal nature. Therefore, these multimodal tokens are as suitable for autoregressive training as textual tokens, allowing us to unify the training objectives for understanding and generation of multimodal tokens into next-token-prediction with cross-entropy loss. The training objective is thus:
@@
where $x_{t}$ represents the discrete multimodal tokens, and $\theta$ denotes the parameters of the LLM backbone. We use the pre-trained model, Yi-6B-Base (AI et al., 2024), for initialization.
Furthermore, to eliminate the computational inefficiency caused by <pad> tokens, we use the masked packing strategy (Lu et al., 2023; Liu et al., 2024; Dehghani et al., 2023). Specifically, the samples are concatenated along the sequence length dimension until the context window is full. Then, we construct the causal attention mask for the tokens of each sample and mask out all the tokens of the other samples.
@@
After the generation of multimodal tokens, it is essential to use modality-specific decoders to reconstruct the images or speech from the codes. Specifically, for image tokens, we directly utilize SEED-Tokenizer’s decoder, which involves an MLP projection to convert the discrete codes into dense latents. These latents condition an off-the-shelf diffusion model (Rombach et al., 2022) to generate the images in the pixel space (Ge et al., 2023a). The vanilla SpeechTokenizer (Zhang et al., 2023b) involves generating timbre tokens through a non-autoregressive model outside the language model, and then feeding the concatenated content and timbre tokens into the SpeechTokenizer decoder to synthesize speech. In our work, to inject the timbre priors into the multimodal language model itself, the timbre tokens are also generated by the autoregressive language model.
@@ 2.2 Pre-Training
As shown in Table 2, we use a three-stage strategy for pre-training, with each stage targeting different objectives. The three stages are: (1) Alignment Pre-training: This stage focuses on learning a multimodal representation more aligned with the language space. (2) Interleaved Pre-training: This stage aims to obtain a multimodal representation with richer contextual semantics. (3) Speech-enhanced Pre-training: This stage specifically enhances the model’s speech-related capabilities, while concurrently replaying data from other modalities. For more details on the pre-training data and its processing procedures, we refer the readers to Appendix A.
Table 2: Pre-training stages and their details. We use “Inter” to denote “Interleaved” for short. We provide batch sizes for each data type per GPU in image-text pair data:language-only data:(image-text interleaved data + video data):speech-text pair data. See Appendix A and Appendix B for more details including pre-training data sources, data cleaning procedures, pre-training hyperparameters, etc.
@@
In this stage, we extend the data used for pre-training to include image-text interleaved data (including video-text data) as a novel image-text dependency pattern. The image-text interleaving pattern has a different nature compared to pairing patterns. Although Li et al. (2023b) and Sun et al. (2023c) argued that interleaved image-text data mainly serves for multimodal in-context learning, we argue that it is also essential for context-aware image generation where images are generated based on specific context, rather than a precise description of the image content. For example, in image-text interleaved data, the text might serve as the image’s preceding or continuing context, rather than its description. This pattern significantly differs from the previous descriptive image generation demonstrated in image-text paired data, where images are generated based on precise and detailed text that clearly describe the content of the images (Team et al., 2023). Therefore, context-aware image generation is essential for tasks such as chain-of-visual-thought reasoning or visual storytelling (Team et al., 2023; Huang et al., 2016), where images are generated without textual descriptions. Due to the lack of benchmarks and evaluation metrics for context-aware image generation, we provide some demonstrations in §3.5 to showcase the potential of our model in visual storytelling, interleaved video-text generation, instructional image editing, chain-of-visual-thought reasoning, multimodal in-context learning, etc.
@@
We refer the reader to Appendix B for more details about the hyperparameters and prompt templates.